In a classic carnival ride, patrons stand agains I need the answers for 5 and 6 parts
Answer #1

Answer to part 5:

Since the rider is standing against the wall, the frictional force is acting upwards and is equal to the weight of the rider, i.e,   F_{fr}= mg. The normal force acting on the rider is given by : F_N= \frac{mv^2}{r}

This acts towards the centre of the cylinder. Now, F_{fr}=\mu F_N\Rightarrow \mu = \frac{g}{v^2r}

Therefore, the coefficient of friction will decrease by increasing the radius of the ride and increasing the speed of the ride.

Answer to part 6:

Since F_N\leq 1.9mg, so the maximum normal force is :

(F_N)_{max}= 1.9mg = \frac{mv_{max}^2}{r}\Rightarrow v_{max}^2=1.9gr

Hence, \mu_{min}=\frac{g}{v_{max}^2r}=\frac{1}{1.9r^2}=0.014 (Ans)